December 15, 2022

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 11:00 (am) via Zoom and in person in the counting house meeting room (2nd floor)

  • Zoom remote connection information:
 Join ZoomGov Meeting
 Meeting ID: 160 710 0885
 Passcode: 595071
 One tap mobile
 +16692545252,,1607100885# US (San Jose)
 +16468287666,,1607100885# US (New York)
 Dial by your location
       +1 669 254 5252 US (San Jose)
       +1 646 828 7666 US (New York)
       +1 669 216 1590 US (San Jose)
       +1 551 285 1373 US
       833 568 8864 US Toll-free
 Meeting ID: 160 710 0885
 Find your local number:

Useful Links

Current Status

  • All shifts have been filled! Thanks Hakob!!!
  • TODAY (12/15)
    • Finished CSB data collection during OWL.
    • Beginning Large Angle, Large Momentum HMS electron settings.
    • BCM calibration began around 10:00 AM
  • YESTERDAY (12/14) oh boy...
    • Attempted Beam Energy Measurement during the morning begging at 9 AM. Did NOT go well.
      • The transport tunnel and the hall was filled with radiation, which caused several issues. CARM strip chart comparison
      • About 7x higher than the measurement taken June 12. Prevented controlled access.
    • At 11:40 AM, heater power and fan powered went to zero. Experts called in.
      • Eventually discovered by Chris Keith and Dave Meekins that the UPS power strip had died (due to radiation).
    • ODH alarms in the transport tunnel and the west wall went off, preventing controlled access. OM4, OM3, both now bypassed. Later discovered to have been ODH HIGH alarms (failure due to radiation).
    • Cryo group was concerned that there was a leak with all the chaos. Cryo was turned off to the targets. False alarm. Cryo returned to the target around 3:30 PM.
    • Targets quickly cooled back down and made a full recovery. No He3 or He4 loss expected.
    • Began to bring back the hall to beam permit at 3:30 PM. Tune beam delivered at 5:30 PM. Smooth running since. Except for a few SHMS runs had the wrong prescales
    • CSB kinematics finished.

Current Run Plan

  • Continue following run plan with Large Angle HMS data at electron polarity.

Notices in Effect for Shift Crew

  • Maximum current on the cryo targets is 40 uA
  • Adjust prescalers as needed to keep SHMS rate under 8 kHz and and HMS under 4 kHz.

Current Issues

  • SHMS pre shower HV channels keep failing. Dave has moved 2 channels around so far. sPre14+, sPre9+
  • HMS Q2 alarm handler doesn't register positron polarity and the limits have to be changed manually.
  • HMS NMR lock struggling. Was locked when it was set, so things are probably fine.

Planned Accesses to the Hall

Planned Accelerator Configuration Changes

Special Requests